H3AsO4 + Mn3P7 + Cr3AsO4 --→ CrP2 + MnAsO4 + H3P 1. Write the oxidation half reaction. 2. Identify the reducing agent in this reaction. 3. Balance the equation. 4. Name the manganese containing product in the equation.
Answer #1

First balance the reaction.

Balanced equation is 11 H3AsO4 + 4 Cr3AsO4 + 5 Mn3P7 = 12 CrP2 + 15 MnAsO4 + 11 H3P

Instructions on balancing chemical equations:

  • Enter an equation of a chemical reaction and click 'Balance'. The answer will appear below
  • Always use the upper case for the first character in the element name and the lower case for the second character. Examples: Fe, Au, Co, Br, C, O, N, F.     Compare: Co - cobalt and CO - carbon monoxide
  • To enter an electron into a chemical equation use {-} or e
  • To enter an ion specify charge after the compound in curly brackets: {+3} or {3+} or {3}.
    Example: Fe{3+} + I{-} = Fe{2+} + I2
  • Substitute immutable groups in chemical compounds to avoid ambiguity.
    For instance equation C6H5C2H5 + O2 = C6H5OH + CO2 + H2O will not be balanced,
    but XC2H5 + O2 = XOH + CO2 + H2O will
  • Compound states [like (s) (aq) or (g)] are not required.
  • If you do not know what products are enter reagents only and click 'Balance'. In many cases a complete equation will be suggested.
  • Reaction stoichiometry could be computed for a balanced equation. Enter either the number of moles or weight for one of the compounds to compute the rest.